OpenStudy (phoenixfire):

How would I show that the error function erf is strictly increasing? I got the Taylor series: \[\sum _{ n=0 }^{ \infty }{ \frac { (-1)^{ n }x^{ 2n+1 } }{ n!(2n+1) } } \] and I applied \[\frac{a_{n+1}}{a_n} > 1\] however it does not work out to be greater than 1. \[\frac { a_{ n+1 } }{ a_{ n } } =\frac { (-1)^{ n+1 }x^{ 2(n+1)+1 } }{ (n+1)!(2(n+1)+1) } \cdot \frac { n!(2n+1) }{ (-1)^{ n }x^{ 2n+1 } } =\frac { (-1)(-1)^{ n }x^{ 2n }x^{ 3 } }{ n!(n+1)(2n+3) } \cdot \frac { n!(2n+1) }{ (-1)^{ n }x^{ 2n } x } \] \[==\frac { (-1)(2n+1) }{ (n+1)(2n+3) }x^{ 2 }\ngtr 1\]

OpenStudy (nipunmalhotra93):

Wait a min. Is the taylor series given? Or did you find it?

OpenStudy (phoenixfire):

I found it, and it's known. Obviously I've ignored the \(\large \frac{2}{\sqrt{\pi}}\) that is multiplied by all of it. http://en.wikipedia.org/wiki/Error_function#Taylor_series

OpenStudy (kainui):

@nipunmalhotra93 Actually it's not too hard to find the taylor series of erf, just plug in (-x^2) for the power series of e^x and then integrate it, that's why you have the (2n+1) in the power and divided by, just the power rule.

OpenStudy (phoenixfire):

I guess if I apply a different property for increasing if \(a_n=f(n),\forall n \in \mathbb{N}\) and \(f'(x) \gt 0, \forall x\in [1,\infty)\) since \(f'(x)=\frac{2}{\sqrt{\pi }}e^{-x^2}\) then as \(x \rightarrow \infty\), \(f'(x) \rightarrow 0\) but never reaching 0 it'll still satisfy the property stated above, and therefore be strictly increasing? Is this a valid argument?

OpenStudy (kainui):

I think you're right, that looks like the way to go since you don't have to worry about infinity; infinity isn't part of your domain.

OpenStudy (phoenixfire):

@Kainui Alright, I'll just stick with that way. And your suggestion looks like it works, it gets rid of the pesky -1 and the x term ends up to the power of 4, so it's all positive and then when multiplied by the ignored factor \(\large \frac{2}{\sqrt{\pi}}\) it's all greater than 1. But I don't like that logic (and too much algebra), sorry lol. I'll go with the derivative method.

OpenStudy (kainui):

Haha yeah fair enough. I honestly don't see why it shouldn't be true. After all, if \[\frac{a_{n+1}}{a_n}>1\] then obviously since n is arbitrary, \[\frac{a_{n+2}}{a_{n+1}}>1\] Just rearranging the inequalities: \[a_{n+2}>a_{n+1}\]\[a_{n+1}>a_{n}\]then\[a_{n+2}>a_{n}\] \[\frac{a_{n+2}}{a_{n}}>1\] It seems to be alright.

OpenStudy (phoenixfire):

But by the same logic of rearranging the inequalities \[\frac{a_{n+1}}{a_n}>1\Rightarrow \frac{a_n}{a_{n+1}}<1\] which is true as it's just the reciprocal of my first attempt which is negative...

OpenStudy (kainui):

Nah, that's not right because dividing by a negative in an inequality reverses the direction of it, so it will still be ok.

OpenStudy (kainui):

You've just abstracted it away into a variable. Plus, it's false as it is, since no negative number is larger than 1. Well maybe. Now I am either confusing you or myself. Whatever. I'm done causing problems now lol.

OpenStudy (phoenixfire):

Hahaha, I'll just leave it as it. Thanks for your help!